OpenStudy (anonymous):

so i have this math problem. it says solve each pair of equations using elimination or substitution. i chose substitution. heres the problem: 2x-3y=1 3x=5y=11 i sort of had an idea of what i was doing and got to 3x= 1-5y divide the 3 and the 5 and the 1 by 3 and you get x=1/3-5/3 next i replaced all the x's with that so it looked like this: 2(1/3-5/3y) - 3y=1 so it woud be 2/3-10/3y-9/3 y= 1 2/3-19/3y=1 now i am stuck i know that i have to subtract 2/3 from both ends and i'll end up with -19/3y=1/3 thats where im at right now

OpenStudy (anonymous):

your secong equation hass two = signs and i'm think one of theme should be a + or a - sign. also after you divided by 3 you said you got x = 1/3 - 5/3 but it should be x = 1/3 - 5y/3.

OpenStudy (anonymous):

hi. ummm as you can probably see i am having some problems. i'm pretty sure i have to retrice-19/3 to 1 but i seriously have no idea

OpenStudy (anonymous):

oh sorry it is a - i am so sorry i'm a cruddy typist:)

OpenStudy (anonymous):

then combine like terms... keep the variables in the numerator unless they started in the denominator or you divided by it.

OpenStudy (anonymous):

oh wait no its a +. yet again sorry:)

OpenStudy (anonymous):

so in your last... 2/3 - 19y/3 = 1 just subtract 2/3 from each side and then multiply both sides by the reciprocal of the coefficient on the y term

OpenStudy (anonymous):

alright i understood almost all those words. whats a coefficient?

OpenStudy (anonymous):

-19y/3 the coefficient is the number that multiplied by y to give you that expression... the coefficient is -19/3

OpenStudy (anonymous):

Coefficients Coefficients are the number part of the terms with variables. In 3x2 + 2y + 7xy + 5, the coefficient of the first term is 3. The coefficient of the second term is 2, and the coefficient of the third term is 7.

OpenStudy (anonymous):

so like it would be 2/3-19/3y=1 -2/3 -2/3 -19/3y= 1-2/3(or 1/3) depending on what you want i guess?

OpenStudy (anonymous):

i mean honestly i have no idea. what do you want me to do and try to use language that someone really stupid would understand please

OpenStudy (anonymous):

i subtracted 2/3 from the front and end. now what do you want me to do. by the way i understand fully what a coefficiant is now. thanks for that wonderful tidbit of info

OpenStudy (anonymous):

no... i want you to learn it. math is a language (believe it or not). you have to learn what the terms mean. what do you like? i'm sure it has it's own "language" too! terms that someone outside wouldn't understand. it happens...

OpenStudy (anonymous):

-19/3y=1-2/3 ( would this equal 1/3 or what)

OpenStudy (anonymous):

so the coefficient is -19/3. do you know what reciprocal means?

OpenStudy (anonymous):

yes, the right hand side is 1/3... very good

OpenStudy (anonymous):

yeah if you have the fraction 8/6 then the recipricol would be 6/8 its a flip flop

OpenStudy (anonymous):

exactly... so what's the reciprocal of -19/3?

OpenStudy (anonymous):

3/-19 easy peasy lemon squeezy

OpenStudy (anonymous):

there you go... multiply both sides of your equation by that number to solve for y

OpenStudy (anonymous):

3/-57 that would simplify to -19 wouldnt it?

OpenStudy (anonymous):

-(3/19)*(19y/3) = -(3/19)*(1/3) 57y/57 = -3/57 y = -1/19

OpenStudy (anonymous):

there's another error though... you used the second equation and solved for x right?

OpenStudy (anonymous):

but don't worry... we'll only have to redo the last few steps

OpenStudy (anonymous):

-19y/3 = 11 - 2/3

OpenStudy (anonymous):

that doesnt make sense because you aid multiply the recipricaol and i did. you said multiply -19/3 by its recipricaol and it dissapppears. the yuo only have x then i multiplied -1/3 by 3/-19 and i got -57/3 or whatever anf that gave me -19. where am i messing up. seriously i am so confused and i dont want to sound like a drama queen but i mean come on. i dont want to "learn the language" i need to know how to do this and youre confusing me and now im all upset and i sound all crazy when really i just want to fuininsh my summer assignmeant. i already hVE THA ANSWER. EVENTUALLY WHATEVER YOURE TRYING TO TELL ME QILL TURN INTO(2,1) but i dont see that happening at this point

OpenStudy (anonymous):

good luck

OpenStudy (anonymous):

ok :) the first step is to isolate either the first or second equation.

OpenStudy (anonymous):

Then either isolate "x" or "y"

OpenStudy (anonymous):

first off thank you. and are we starting from the beginning or where?

OpenStudy (anonymous):

Yea lets start from he beginning because I also have to get familiar with the problem

OpenStudy (anonymous):

So pick an equation and then pick a variable

OpenStudy (anonymous):

k well i subtracted 5y from 11 leaving me with 3x=11-5y am i ok thus far?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Keep in mind that you can also choose to isolate y if it looks easier

OpenStudy (anonymous):

alright then i divided everything by 3 leaving me with x= 11/3-5/3y all good?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

then i brought all that into the other equation 2x-3y=1 giving me 2(11/3-5/3y)-3y=1 is this ok

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

this would give me 22/3-10/3y-3y=1 this would give me 22/3--10/3-9/3 that would eventually give me 22/3-1/3=1 yes?

OpenStudy (anonymous):

forgot the y

OpenStudy (anonymous):

so correct?